OpenStudy (anonymous):

can someone show me how to use long division to rewrite (x^2-3x+2)/(x+1) and is there any way of knowing if you need to use long division on an integral?

terenzreignz (terenzreignz):

Attempting long division on integrals, more likely than not, will make it harder :)

jhonyy9 (jhonyy9):

so you see that x^2 -3x +2 can being factorized easy like (x-2)(x-1) - and now because there not is like factor (x+1) so hence this signing that you need using long division

OpenStudy (anonymous):

thx johnny and terenz long division is sometimes the best option

terenzreignz (terenzreignz):

but if the denominator is just a linear function of the form ax + b, and the numerator is another polynomial, then I'd just use integration by substitution, that way, no tedious long division (I'm not good at it :) )

OpenStudy (anonymous):

yeah but that doesnt always work like take the integral of the function i mentioned and youll see that u substitution and factoring dont get you anywhere

terenzreignz (terenzreignz):

I'll try :)

terenzreignz (terenzreignz):

\[\int\limits_{}^{}\frac{x^{2}-3x+2}{x+1}dx\]

terenzreignz (terenzreignz):

I let u = x + 1 du = dx \[\int\limits_{}^{}\frac{(u-1)^{2}-3(u-1)+2}{u}du\] \[\int\limits_{}^{}\frac{u^{2}-2u+1-3u+3+2}{u}du = \int\limits_{}^{}\frac{u^{2}-5u+6}{u}du\] \[\int\limits_{}^{}(u - 5 + \frac{6}{u})du= \frac{u^{2}}{2}-5u+6\ln|u| + C\] \[\frac{(x+1)^{2}}{2}-5(x + 1)+6\ln|x+1| + C\] How's that? :)

OpenStudy (anonymous):

that is a very complicated u substitution and in the time it took u to do it i taught myself long division and solved the integral

terenzreignz (terenzreignz):

To be fair, what took a long time was typing it up

OpenStudy (anonymous):

still though this would be the most complicated u sub ive done so far and this long division isnt as hard as i thought it would be before

terenzreignz (terenzreignz):

It doesn't matter, as long as the process is legit and you arrive at the same correct answer- and besides, you said I couldn't do it by u substitution ;)

OpenStudy (anonymous):

and the answer in the book and my answer came out to (x^2)/2 -4x +6ln(x+1)+C which is probably the same answer as yours just more simplified

OpenStudy (anonymous):

and i didnt forget the absolute bars i just cant do em on here

terenzreignz (terenzreignz):

It's the same, lol, and yeah, more simplified, I didn't feel like simplifying further :D

OpenStudy (anonymous):

soo i arrived at my answer faster and more simplified which is crucial on tests so now i have another method in my arsenal to tackle integrals!

terenzreignz (terenzreignz):

I'd simplify it if it was on paper, though. And uhh, you being faster is debatable... :)

OpenStudy (anonymous):

well if i did it your way it would have taken ME longer but seeing it done your way will help me if i forget how to do long division so thx :)

terenzreignz (terenzreignz):

uhh, sure no prob... (btw, I never imagined using long division for integrals, just a pet peeve, I guess)

OpenStudy (anonymous):

i never imagined using long division past elementary school :P